Assume that you own 110 shares of common stock of a company, that you have been receiving cash dividends of $6 per share per year, and that the company has a 5-for-4 stock split.

Required:

  1. How many shares of common stock will you own after the stock split?
  2. What new cash dividend per share amount will result in the same total dividend income as you received before the stock split? (Do not round intermediate calculations. Round your answer to 2 decimal places.)
  3. What stock dividend percentage could have accomplished the same end result as the stock split?
Answer #1

Solution:

Nos of shares after stock split = 110*5/4 = 137.50 shares

New cash dividend per share = (110*$6) / 137.50 = $4.80 per share

Stock dividend percentage that accomplished the same end result = (137.50 - 110) / 110 = 25%